Required information The following information applies to the questions displayed below.] Sandra would like to organize BAL as either an LLC (taxed as a sole proprietorship) or a C corporation. In either form, the entity is expected to generate an 13 percent annual before-tax return on a $750,000 investment. Sandras marginal income tax rate is 37 percent and her tax rate on dividends and capital gains is 23.8 percent (including the 3.8 percent net investment income tax). If Sandra organizes BAL as an LLC, she will be required to pay an additional 2.9 percent for self- employment tax and an additional 0.9 percent for the additional Medicare tax. BALs income is not qualified business income (QBI) so Sandra is not allowed to claim the QBl deduction. Assume that BAL will distribute all of its after-tax earnings every year as a dividend if it is formed as a C corporation. (Round your intermediate computations to the nearest whole dollar amount.) a. How much cash after taxes would Sandra receive from her investment in the first year if BAL is organized as either an LLC or a C corporation? After-tax cash flow LLC C corporation 58,693
Answer #1

a. Calculation of cash after taxes if BAL is organised as LLC

Return = 750000*13% = $ 97500

Less - Income tax @ 37% = $36075

Self employment and medicare tax = $ 3705

After tax cash flow = $57720

b. Calculation of cash after taxes if BAL is organised as C corporation

Return = 750000*13% = $ 97500

Less - Income tax @ 37% = $36075

Tax on dividents @ 23.8% = $ 14619

After tax cash flow = $ 46806